OpenStudy (anonymous):

HELP PLEASE!!! Solve the equations and check your solution, 2/5+1/y-5=y-7/5y-25

OpenStudy (anonymous):

\[is \it \frac{ 2 }{5 }+\frac{ 1 }{y-5 }=\frac{ y-7 }{5y-25 }\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 2\left( y-5 \right)+5 }{5\left( y-5 \right) }=\frac{ y-7 }{5y-25 }\] \[\frac{ 2y-10+5 }{5y-25 }=\frac{ y-7 }{ 5y-25 },or \frac{ 2y-5 }{ 5y-25 }=\frac{ y-7 }{ 5y-25 }\] now you can solve.